OpenStudy (anonymous):

help please!:) Expand the series and evaluate: Enter your answer as the example: 1+2+3+4+5=15

OpenStudy (anonymous):

\[\sum_{k=1}^{5} [8(k+1)]\]

OpenStudy (anonymous):

16+24+32+40+48=160

OpenStudy (anonymous):

\[\sum_{k=1}^{5}\left[ 8\left( k+1 \right) \right]\] =8(1+1)+8(2+1)+8(3+1)+8(4+1)+8(5+1) =8(2+3+4+5+6) =8*20=160

OpenStudy (anonymous):

thanks guys!:)